Factor as much as possible x^2 + x - 6
Answer by jorel1380(2518)

x^2 + x - 6 = (x+3)(x-2)
x^2 + x - 6 = x^2 + 3x - 2x - 6
x^2 + x - 6 = x^2 + x - 6 = (x+3)(x-2)